

********************************************************************
UARS WEEKLY STATUS REPORT - 25 MAR 2005
********************************************************************
***********
SOLSTICE
***********
SOLSTICE began day only operations on March 21 in support of the yaw
maneuver. Otherwise, the instrument continues to perform exceptionally,
taking solar data on all available orbital tracking periods. An average
of 45% of data were returned this week accounting for 267 minutes of
solar observing time.
***********
SUSIM
***********
SUSIM was turned off after measurements on March 20. Earlier, SUSIM
had successfully gathered daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring scans
for time periods with UARS telemetry or tape recording.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 12 March 2005 (Orbit 73887, GMT Day 071) through
18 March 2005 (Orbit 73993, GMT Day 077).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 073/1348 73926 POWER OFF
HRDI 076/1415 73972 POWER ON - DAY ONLY
The beta angle ranged between the angles of 35.6 degrees to 14.4
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 3.0 degrees on 21 March 2005 (DOY 080).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 3.65 to 2.24 degrees C. Additional
battery performance data for each battery on 11 March 2005 and 18
March 2005 are provided in Appendix D.
The clock error ranged from 11.5 to -10.0 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 02
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 05
HALOE = 01
SUSIM = 01
HRDI = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 08
Five (5) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Six (6) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to monitor the batteries closely. Early
on 14 March 2005 the max temperature of battery 2 climbed to 12.64 C
and the temperature delta between batteries 2 and 3 reached 5.44 C.
Due to these high values the FOT powered of the HALOE instrument and
commanded VT-6. In the reduced VT level both the battery 2 temp and
temperature delta came down within a few orbits. This charge level
was maintained until late 17 March 2005 when the battery 3
differential voltage dropped to -324 mV. VT-7 was then commanded to
correct this problem. The MPS continued to utilize the VT-7 charge
rate for the remainder of the week.
The FOT has completed their spacecraft tests with loads built on the
new CMS software. All tests results were positive and the official
release will be installed on the prime CMS on 21 March 2005. When the
install is complete, UARS will have the capability to properly
schedule SMA events. UARS will begin using this functionality
beginning 28 March 2005. A decision from the UARS project scientist
is needed to determine the frequency of its use given the current
battery issues.
MEETINGS
The FOT attended the biweekly UARS clock rollover meeting on 01 March
2005 to the timeline of activities and current action items to
complete this task.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 21 March 2005 (DOY 080). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 18 March 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
73893 071/0852 05-067 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
73926 073/1348 05-068 HALOE POWER OFF / CHANGE VT TO L6
73930 073/2043 05-069 PMON PMON 31 CLEAR/ACTION 21 ENABLE
73942 074/1534 05-070 SUSIM SUSIM COMMANDS
73968 076/0755 05-071 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
73972 076/1415 05-072 HRDI POWER ON - DAY ONLY
73978 077/0108 05-073 MPS VT LEVEL 7
73986 077/1353 05-074 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 18 March 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
OBSPMGPOOL31 PMON GROUP 31 OOL 73897
OBSPMACOOL21 PMON ACTION 21 OOL 73897
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 18 March 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
73893 071/0852 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
73968 076/0755 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
73986 077/1353 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 18 March 2005
11 March 2005 - GMT Day 070 Beta = 43.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -1.80/-2.74 +00.0/+00.0 07.0 0.00 0.0
2 218.4/ 78.4 +11.76/+10.40 +26.8/-15.2 26.7 1.10 19.7
3 -33.6/-134.4 +7.41/ +6.43 +24.0/-13.2 26.7 1.06 18.2
18 March 2005 - GMT Day 077 Beta = 14.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.42/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 336.0/ 84.0 +10.74/+7.09 +24.0/-12.8 27.0 1.10 18.6
3 -33.6/-140.0 +6.11/+3.87 +21.6/-11.2 27.0 1.07 16.9
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
73926 073/1348 LV7 -> LV6
73978 077/0108 LV6 -> LV7
APPENDIX E
UARS ANOMALY REPORTS
Week ending 18 March 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-012 FOT 73894 071/1105 30832 Generic Late Acquisition
05-015 MPS 73895 071/1606 N/A PMON grp 31/PMON act 21
trip
05-012 FOT 73924 073/0945 30832 Generic Late Acquisition
05-016 FOT 73926 073/1348 N/A Data dropout; bad
weather@Guam
05-017 FOT 73926 073/1423 38922 Data dropout; IP-NOC
linework
05-018 FOT 73930 073/2043 N/A PMON grp 31/PMON act 22
trip
05-012 FOT 73976 076/2108 30832 Generic Late Acquisition
05-012 FOT 73989 077/1805 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-013 FOT 73876 070/0549 38897 Drop Out on the Return
Link
05-014 FOT 73881 070/1351 30824 Data Dropout
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-012 FOT See List Mar 05 30832 Generic Late Acquisition
05-015 MPS 73895 071/1606 N/A PMON grp 31/PMON act 21
trip
05-016 FOT 73926 073/1348 N/A Data dropout; bad
weather@Guam
05-017 FOT 73926 073/1423 38922 Data dropout; IP-NOC
linework
05-018 FOT 73930 073/2043 N/A PMON grp 31/PMON act 22
trip